A(1,1) and B(2,-3) are two points and D is a point on AB produced such that AD=3AB. Find the co-ordinates of D

Grade:11

3 Answers

AJIT AskiitiansExpert-IITD
13 years ago

Dear Chanchal ,

we are given AD =3AB , AB +BD  = 3AB ;  2AB = BD , therefore B divides AD in the ratio of 1:2.

let d be (x,y ) then applying formulae ,

   2  = (x +2)/3 ;  x = 4 and -3 = (y +2)/3 =-3 ; y =-11

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Win exciting gifts by answering the questions on Discussion Forum.


gOlU g3n|[0]uS
13 years ago

ans    let D(x,y)

           let AB=k and AD= 3k

           x=3k(2)+k(1)/4k = 6k+k/4k = 7/4.

           y=3k(-3)+k(1)/4k = -9k+k/4k = -2.

   HENCE D(7/4,-2).Smile

Madesh
5 years ago
AD = 3AB
SO, AB +BD = 3AB
BD = 2AB
Therefore, using formula, 
 2 = (X+2)/3 = x ; and -3 = (y+2)/3 = y
So, X= 4 and y = -11